OpenStudy (lovelyharmonics):

If the formula for a compound is represented by X2Y3 and the charge on the Y ion is -2, what is the charge in the X ion?

OpenStudy (lovelyharmonics):

@ikram002p

OpenStudy (ikram002p):

(x) × 2+(-2 ) × 3=0 solve for x

OpenStudy (lovelyharmonics):

-2*3=-6 2-6=-4

OpenStudy (ikram002p):

mm sorry -4 is wrong 2 x -6=0 2x=6 x=6/2=3